OpenStudy (anonymous):

At the Ski Center, the most people will pay for a neck warmer is $10. The manager of the Ski Center knows he needs a 30% markup based on cost. What is the most the manager can pay the suppliers for neck warmers and still keep the selling price at $10?

OpenStudy (anonymous):

a 'markup' price equation will look like this: $10 = SupplierPrice + SupplierPrice*30% so it can be written like this: $10 = P + P*30% $10 = x + x*30% $10 = 1.0*x + 0.3*x then solve for, 'x' ^_^

OpenStudy (anonymous):

ok thank you ;-)

OpenStudy (anonymous):

ur welcome, the idea will be identical in solving other 'markup' problems ^_^